OpenStudy (anonymous):

Name the binomial you can multiply by (x + 9) to get the product x^2+5x-36 A. 4 - x B. x + 6 C. x - 4 D. x - 6

OpenStudy (hoblos):

(x + 9)(x - 4) = x^2 -4x+9x -36 = x^2+5x-36 so C